OpenStudy (anonymous):

Simplify. √25x^6 times √36y^4

OpenStudy (anonymous):

Is this your question? \[\sqrt{25x^6} \times \sqrt{36y^4}\]

OpenStudy (anonymous):

yes

OpenStudy (goformit100):